OpenStudy (anonymous):

Write the number 2.3171717... as a ratio of integers

OpenStudy (anonymous):

x = 2.31717... 1000x = 2317.1717... 1000x-x = 2317 999x = 2317 x = 2317/999 That ratio might be simplified, idk

OpenStudy (anonymous):

wrong.

OpenStudy (anonymous):

the method is right though...

OpenStudy (anonymous):

ahaha you're right. I can't even do simple arithmetic

OpenStudy (anonymous):

try subtracting 10x :-D

OpenStudy (anonymous):

x = 2.31717... 1000x = 2317.1717... 1000x-x = 2315 999x = 2315 x = 2315/999

OpenStudy (anonymous):

even 2315/999 is wrong :-(

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

x = 2.31717... 1000x = 2317.1717... 10x = 23.1717... 1000x-10x = 2294 990x = 2294 x = 2294/990 lol

OpenStudy (anonymous):

yay that's the right answer. thanks